Kendra moves a wagon weighting 1500 grams along a flat road by pulling with a force of 10 newtons with an angle of 37 degrees above the horizontal. If the wagon started at rest and that friction is negligible, how much work does she do on the wagon in 4 seconds?


1
Expert's answer
2020-07-24T10:08:34-0400

Acceleration:"a = \tfrac{F}{m}" "\\tfrac{10cos(37^{\\omicron})}{1.5}\t= 5.32 ( \\tfrac{m}{s^{2}})"

Distance:"d= \tfrac{at^{2}}{2}"

"d = \\tfrac{5.32*4^{2}}{2}\t = 42.6m"


Work:

"W=F*d"


"W = 10cos(37^{\omicron})*42.6 = 340 J"


Final Answer:

"Work = 340 J"